Under which mathematical principle is created the Newton Interpolating polynomial

a

Lagrange polynomial

b

Divided Difference technique

c

Limit definition

d

Subtracted difference technique
Answer #1

Ans: (b) divided difference techniques

Why?:

Newton interpolating polynomial passing through k+1 points is given as

η j-1 Ρ, (x) = f(x)+f[xo,X1, 2, ...,x] (Πα - (x – x) J=1 =0

Where is called kth divided difference and it’s formula is

-2 f(x;) f[x0, x1,x2,......,x;]= -Em-julk: - x;) for all k = 0,1,2,3,..... ....n

So newton interpolating polynomial is created based on divided difference technique.